OpenStudy (moongazer):

let x=4 -x^4 = -256 is this correct?

OpenStudy (moongazer):

not -x^4 = 256 because this -x^4 means -1(4)^4 not (-4)^4

OpenStudy (moongazer):

Is it correct?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

it is 256

OpenStudy (theviper):

(-x)^4=256 & -(x^4)=-256

OpenStudy (anonymous):

becoz instead of x u r giving 4

OpenStudy (theviper):

Does that help @moongazer ??

OpenStudy (anonymous):

-4 * -4 * -4 * -4=

OpenStudy (theviper):

Becoz according to BODMAS rule first we solve brackets :)

OpenStudy (moongazer):

so -x^4 = -256 is correct right?

OpenStudy (anonymous):

but it is like this (-4)^4

OpenStudy (moongazer):

But there is no parenthesis ?

OpenStudy (moongazer):

so it is like -1(x)^4

OpenStudy (anonymous):

or wat is -x*-x*-x*-x

OpenStudy (anonymous):

nw give x=4

OpenStudy (anonymous):

wat will the answer

OpenStudy (anonymous):

just think like that

OpenStudy (anonymous):

-x^4=-x*-x*-x*-x

OpenStudy (moongazer):

isn't that \[-4^2 \neq (-4)^2\]

OpenStudy (moongazer):

???

OpenStudy (anonymous):

why both r same

OpenStudy (moongazer):

no it is not

OpenStudy (anonymous):

why>>>>>>

OpenStudy (moongazer):

http://www.wolframalpha.com/input/?i=-4%5E2

OpenStudy (anonymous):

ok.....@moongazer nw tell is this correct or not -x^4=-x*-x*-x*-x

OpenStudy (moongazer):

http://www.wolframalpha.com/input/?i=%28-4%29%5E2

OpenStudy (anonymous):

@moongazer

OpenStudy (moongazer):

That is wrong I think.

OpenStudy (lgbasallote):

-x^4 = -256 for the record...

OpenStudy (anonymous):

indeed

OpenStudy (moongazer):

that's the question I am asking :)

OpenStudy (lgbasallote):

so i guess it's down to a 2-on-2

OpenStudy (anonymous):

why //

OpenStudy (anonymous):

@lgbasallote -x^4=-x*-x*-x*-x

OpenStudy (anonymous):

@Yahoo! this is -(x^4) not (-x)^4

OpenStudy (lgbasallote):

because the exponent just affects the nearest base

OpenStudy (anonymous):

look his question it is just -x^4

OpenStudy (lgbasallote):

exactly

OpenStudy (lgbasallote):

the ^4 is just on x...not on - too

OpenStudy (moongazer):

because -x^4 is like -1(x)^4 not (-x)^4 I think this is the explanation?

OpenStudy (anonymous):

i would say if (-x)^4=256 & -(x^4)=-256

OpenStudy (lgbasallote):

-x^4 = -(x^4)

OpenStudy (anonymous):

so wat is -2^2

OpenStudy (lgbasallote):

-4

OpenStudy (moongazer):

@Yahoo! (-1)(4) so -4

OpenStudy (anonymous):

-2^2 = -1 (2^2) = -1 * 4 = -4

OpenStudy (lgbasallote):

let this be an end to these skepticisms http://www.wolframalpha.com/input/?i=-4%5E4

OpenStudy (anonymous):

when there is a minus, do all the stuff before minusing

OpenStudy (anonymous):

hey ..... it is like this (-x)^4=256 & -(x^4)=-256

OpenStudy (anonymous):

nop visit http://www.wolframalpha.com/input/?i=%28-4%29^4

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-2^2

OpenStudy (anonymous):

(-x)^4=256 -x^4 = -256 ................ agreed

OpenStudy (lgbasallote):

lol